OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1)

OpenStudy (zzr0ck3r):

x-2y = 14 -x-3y=-9 add them -5y = 5 y=-1 x-2(-1) = 14 x+2 = 14 x = 12 (12,-1)